OpenStudy (anonymous):

find dy/dx for ln(5y)=e^y sin(5x)

OpenStudy (anonymous):

hi do you have the selected answers?

OpenStudy (anonymous):

there are many possible forms of answers lol....

OpenStudy (anonymous):

find dy/dx for ln(5y)=e^y sin(5x) Dln(5y)=D[e^y sin(5x)] 5y'/5y =e^y(5cos5x)+(sin5x)y'e^y y'/y-(sin5x)y'e^y=e^y(5cos5x) factoring out y' we get y'(1/y-(sin5x)e^y)=5e^y(cos5x) y'=5e^y(cos5x) /[(1/y)-(sin5x)e^y)] or y'=5/[(1/ye^ycos5x-tan5x)]